OpenStudy (anonymous):

Graph the circle (x + 2)2 + (y + 1)2 = 32. Label the center and at least four points on the circle.

OpenStudy (campbell_st):

well the general from is \[(x - h)^2 + (y - k)^2 = r^2\] (h, k) is the centre and r is the radius.. by comparing your equation, whats the centre of the circle.. and whats the radius..?

OpenStudy (anonymous):

The center will be -2,-1 but what do i do with the 32?

OpenStudy (campbell_st):

well you are looking at the radius as \[r = \sqrt{32} ...or.... r = 4\sqrt{2} \] so the 4 points I'd plot on the circumference are \[x = -2\pm 4\sqrt{2} ...y = -1\] and \[x = -2, ..y = -1 \pm 4\sqrt{2} \] here is what I mean  you need the other 2 points. and it is a strange radius value.